# Electron-Phonon interaction and lattice thermal conductivity from metals to 2D Dirac crystals: a review

### v1

@@ 4.2 The role of different phonon branches
In covalent semiconductors such as Si and Ge, long-wavelength acoustic (LA/TA) modes couple to carriers through the (short-range) deformation-potential interaction. The corresponding e–ph matrix elements vanish at $q\!\to\!0$ and increase with $|q|$; together with the linear acoustic dispersion $\omega=v_{s}q$, this makes carrier scattering by the very lowest-frequency LA/TA modes weak. As a result, these modes retain long mean free paths and dominate the lattice thermal conductivity at room temperature.[60, 62, 63]
In polar III–V semiconductors the situation changes: the long-range Fröhlich field couples carriers strongly to longitudinal-optical (LO) phonons, and the mode dependence of the e–ph matrix elements becomes essential.[61, 64, 65] When electrons and phonons are treated self-consistently in coupled BTE frameworks, momentum/energy exchange mediated by e–ph scattering can noticeably reduce the phonon thermal conductivity relative to a phonon–phonon-only model, especially at finite carrier densities.[66, 67]
Two-dimensional semiconductors add a quadratic flexural (ZA) branch. DFPT together with anharmonic renormalization shows that, at room temperature and for carrier densities $\lesssim 10^{12}\,\mathrm{cm^{-2}}$, ZA scattering contributes less than $\sim\!30\%$ of the total resistive e–ph rate; in-plane acoustic modes thus dominate both carrier mobility and lattice heat conduction under typical conditions.[68]
@@ 4.3 Carrier density and screening
where $D_{\mathrm{ac}}$ is the acoustic deformation potential, $\rho$ the mass density, $v_{s}$ the sound velocity, and $C\sim\mathcal{O}(1)$ is an angular factor that depends weakly on band anisotropy and degeneracy[44, 62].
Second, the dense electron gas strengthens electrostatic screening. Rather than equating the screening vector to $k_{F}$, one should use a Thomas–Fermi/Lindhard wavevector $k_{s}$ that increases with $n$ (via $k_{F}$) but is not equal to it; transport coefficients are notably sensitive to the choice of $k_{s}$[69]. In ab initio BTE studies of $n$-type Si, once $n\!\gtrsim\!10^{19}\,\mathrm{cm^{-3}}$ the calculated electron-limited damping of low-frequency LA/TA modes grows rapidly and the resulting suppression of $\kappa_{\mathrm{p}h}$ tracks time-domain thermoreflectance benchmarks[62, 70].
@@
so increasing $k_{s}$ (heavier doping) softens the long-range field. Nonetheless, room-temperature energy relaxation and phonon damping remain LO-dominated over wide doping ranges in nitrides; screening reduces the rate but does not eliminate the LO bottleneck[71, 72]. In short, heavy doping in covalent Si throttles heat flow primarily by empowering low-frequency acoustics, whereas in polar GaN it leaves the LO channel as the primary sink—highlighting how bonding character (deformation potential vs. Fröhlich) and carrier screening jointly decide which branch ultimately limits the lattice thermal conductivity.
@@ 4.4 Dimensionality and dielectric environment
Reducing a semiconductor to monolayer or ultrathin-film thickness fundamentally modifies the microscopic ingredients of the e–ph interaction. Confinement quantises the out-of-plane phonon wave vector, replacing the 3D continuum $q_{z}$ by discrete values $q_{z}=n\pi/L$ ($n=0,1,\dots$), while the in-plane component $\mathbf{q}_{\parallel}$ remains continuous down to $\mathbf{q}_{\parallel}\!\to\!0$. The in-plane acoustic branch with $n=0$ remains gapless with $\omega=v_{s}q_{\parallel}$; by contrast, the first thickness-quantised subband ($n=1$) acquires a gaplike term,
@@
so long-wavelength 3D channels that require finite $q_{z}$ below $\pi/L$ are kinematically excluded even though arbitrarily small in-plane momenta are allowed. A simple way to encode this thickness cutoff in the out-of-plane phase space is
@@
This $\Theta$-filter is a heuristic to encode the loss of continuous $q_{z}$ phase space; in real films the spectrum consists of discrete subbands (Lamb modes) rather than a hard cutoff. At the same time, freestanding monolayers host a quadratic flexural (ZA) branch with $\omega_{\mathrm{ZA}}=Aq^{2}$. For crystals with horizontal mirror symmetry, the leading e–ph coupling to ZA modes is symmetry-forbidden at first-order and emerges at higher order, giving a small-$q$ scaling $\lvert g_{\mathrm{ZA}}\rvert\!\propto\!q^{2}$; as a result, the intrinsic ZA contribution to resistive scattering is strongly suppressed compared to in-plane acoustics unless symmetry is broken by substrates, strain, or gating.[73] First-principles transport studies in 2D semiconductors consistently find that in-plane acoustic modes dominate and that the role of flexural modes is reduced to $\lesssim\!30\%$ at low carrier densities (with further suppression on substrates).[73, 74]
Confinement also shortens relevant mean-free paths compared with bulk. In monolayer MoS2, electronic mean free paths at 300 K are typically below $\sim\!9$ nm when phonon scattering limits transport[62]. This has been shown in Fig. 3(a). phonon MFPs likewise contract relative to the bulk parents when boundary, substrate, and e–ph mechanisms are included, though precise values are material- and environment-dependent.
A second confinement effect is electrostatic: the long-range Fröhlich interaction in 2D depends on an effective out-of-plane permittivity set by the surrounding media,
@@
where $q_{\text{TF}}$ is the 2D Thomas–Fermi wave vector and $A$ the monolayer area. Increasing $\epsilon_{\mathrm{eff}}$ (e.g. by h-BN encapsulation) weakens the 2D Fröhlich vertex and lengthens LO lifetimes; conversely, high-$\kappa$ substrates (e.g. HfO2) enhance polar damping. First-principles treatments of polar e–ph coupling explicitly include the long-range Fröhlich part of the vertex and its screening,[48, 75, 76] and coupled e–ph BTE solvers now propagate these interactions self-consistently through the electron and phonon distributions under the same driving fields.[70]
@@ 4.5 Coupled BTE and the role of elphbolt
The prevailing ab initio workflow treats electrons and phonons in parallel: one solves a phonon Boltzmann transport equation (BTE) for $\kappa_{\mathrm{p}h}$ and an electron BTE for $\sigma$ and $S$. In reality the two distribution functions are coupled: non-equilibrium electrons push phonons out of equilibrium (electron drag), and drifting phonons pull electrons along (phonon drag). Historically, semi-analytical drag models partially decoupled the BTEs and introduced effective terms, but this can violate the Kelvin–Onsager relationship connecting Seebeck and Peltier responses[66, 77, 78, 79, 80].
@@
elphbolt solves the fully coupled, linearised electron–phonon BTEs from first principles. It iterates the electron and phonon collision integrals to self-consistency under the same driving fields ($\nabla T$ and $E$), and explicitly enforces the Kelvin–Onsager reciprocity at each iteration. Crucially, it uses exactly the same DFPT–Wannier electron–phonon matrix elements as single-species BTE workflows, so one can switch from the independent-channel picture to the coupled one without refitting any parameters[48, 66, 81, 82].
@@
Concrete results illustrate when coupling matters. In $n$-type Si, elphbolt reproduces the temperature dependence of the Seebeck coefficient across low and high doping; the phonon-drag contribution becomes dominant below $\sim\!175$ K, while at room temperature direct e–ph scattering dominates and single-BTE and coupled-BTE predictions nearly coincide. For $\kappa_{\mathrm{p}h}$ in Si, the electron drag correction is small; suppression of $\kappa_{\mathrm{p}h}$ at low temperature is governed mainly by the phonon–electron scattering rates rather than mutual drag. Similar “small drag on $\kappa_{\mathrm{p}h}$” conclusions hold for GaAs and SiC[66, 83, 84]. This has been shown in Fig. 3(b). In strongly polar materials (e.g. GaN), accurate treatment of long-range polar couplings may require quadrupolar corrections; until those are included, drag conclusions should be drawn with care[66, 81, 82].
@@
In short: use coupled e–ph BTEs when phonon or electron drag is relevant (low-to-intermediate $T$, clean samples, or structures with suppressed impurity scattering), or when enforcing Kelvin–Onsager consistency matters for quantitative thermoelectric predictions. Otherwise, single-BTE solutions (RTA or iterative) are often sufficient and substantially cheaper[66].
@@
Figure 3: (a) In monolayer MoS2 at 300 K, phonon-limited electronic mean free paths cluster in the few–10 nm range (often <10<\!10 nm), illustrating the strong size/confinement sensitivity and providing context for the 2D transport lengths discussed in this subsection. Figure taken from Li et al.[62]. (b) Lattice thermal conductivity of Si vs temperature at low and high doping from the fully coupled e–ph BTE agrees with experiment Inyushkin et al.[85] and shows stronger low-$T$ suppression from e-ph scattering. At room temperature ph–ph scattering dominates and coupled/uncoupled predictions converge. Figure taken from Protik et al.[66]
@@
A consistent body of measurements now benchmarks the fully first-principles workflow (DFPT/Wannier e–ph vertices + BTE solvers). In boron-doped silicon, thermal-conductivity data on both natural and isotopically enriched crystals, together with doping-dependent measurements, exhibit the expected suppression of $\kappa_{\mathrm{p}h}$ as carriers are added and temperature is lowered. Fully coupled e–ph BTE calculations (elphbolt) reproduce the measured temperature dependence of $\kappa_{\mathrm{p}h}$ for $n$-type Si across dilute and degenerate limits, and quantify why the phonon–electron channel suppresses $\kappa_{\mathrm{p}h}$ more strongly at low $T$ (where long-wavelength acoustics dominate) than at high $T$ (where ph–ph scattering wins)[66]. These comparisons are made directly against high-purity Si data and are consistent with isotope-controlled measurements (which isolate mass-disorder effects) and with earlier doping-dependent thin-film studies[86, 87]. [^1] In elphbolt’s Si case study, the measured points for $\kappa_{\mathrm{p}h}(T)$ at low and high doping (including $n\!\sim\!2\times 10^{19}\,$cm-3) are reproduced without empirical lifetimes, and electron-drag corrections to $\kappa_{\mathrm{p}h}$ are found to be small; see Fig. 6 and discussion in Ref. [66].
@@
For polar III–V semiconductors (e.g. GaAs), experiments using transient thermoreflectance and related pump–probe probes report reductions of lattice thermal conductivity compatible with LO-phonon damping at finite carrier densities. First-principles treatments that include the long-range polar (Fröhlich) vertex capture these trends; coupled e–ph BTE solutions further show that, while phonon drag can dominate the Seebeck response at low $T$, the electron-drag correction to $\kappa_{\mathrm{p}h}$ remains small in GaAs, consistent with the mode-resolved balance of ph–e versus ph–ph scattering[66, 64].
@@
In 2D semiconductors, measurements on monolayer MoS2 consistently find a strong substrate/encapsulation dependence of the in-plane thermal conductivity, with reported room-temperature values for suspended or weakly coupled monolayers typically tens to $\sim\!$80 W m-1 K-1 and lower values on amorphous oxides[88, 89]. These trends are aligned with first-principles expectations: the dielectric environment tunes the long-range polar coupling, and the dominant in-plane acoustic branches set the heat flux; ab initio transport in monolayer MoS2 also finds short electronic mean free paths and comparatively modest Fröhlich coupling in the intrinsic limit[62]. Encapsulation in high-quality h-BN further improves interfacial heat dissipation and can raise the effective thermal conductance of hBN/MoS2/hBN stacks[90].
@@
Taken together—covalent Si (with isotope/doping controls), polar GaAs (LO-damping trends), and substrate-tunable 2D MoS2—these benchmarks show that once DFPT e–ph matrix elements are in hand, first-principles transport reproduces both magnitudes and temperature trends of $\kappa_{\mathrm{p}h}$ without empirical lifetimes, while clarifying where mutual e–ph drag matters and where it does not[66].
@@ 4.6 Summary and open challenges
The next frontier is to widen this ab initio umbrella. First, correlated or magnetic semiconductors call for beyond-DFT band structures (DFT+U+U, hybrid functionals, or even GWGW/DMFT) so that the subsequent e–ph couplings rest on accurate quasiparticle states.[48] Second, van der Waals heterostructures demand a truly nonlocal description of screening: the effective Fröhlich vertex in, e.g., MoS2/h-BN/graphene stacks depends on the full dielectric profile rather than a single scalar constant, and first-principles long-range corrections must be retained explicitly.[61, 81] Third, pump–probe spectroscopy pushes materials into ultrafast, far-from-equilibrium regimes where hot carriers and nonthermal phonons coevolve; coupling the self-consistent electron– and phonon–BTE engine (elphbolt) to optical drive/relaxation protocols promises a parameter-free view of picosecond energy flow and drag phenomena.[55, 66]
@@
Dirac crystals are gapless systems whose low–energy carriers follow a linear (Dirac) dispersion rather than the usual parabolic one. In graphene, conduction and valence cones touch at discrete Dirac points, and carriers near these points are well described by a massless Dirac Hamiltonian with an approximately constant Fermi velocity $v_{F}$.[91, 92] This “relativistic” band structure underlies characteristic transport phenomena such as the unconventional (half-integer) quantum Hall effect and Klein tunnelling.[91, 93]
@@
In pristine, charge-neutral graphene the Fermi “surface” collapses to a point and the density of states vanishes linearly at the Dirac point. Modest electrostatic gating shifts $E_{\mathrm{F}}$ by tens to hundreds of meV, and—since the 2D Dirac DOS grows $\propto|E|$—yields large changes in carrier density.[91, 92] Doping then affects electron–phonon (e–ph) physics in two ways: it enlarges the electronic phase space for scattering and strengthens screening. The net impact is mode- and $\mathbf{q}$-selective: short-range deformation-potential couplings are efficiently screened, whereas the “gauge-field” part of the in-plane acoustic coupling remains essentially unscreened; moreover, in suspended graphene the linear coupling of out-of-plane (ZA) flexural modes is symmetry-forbidden at first-order and arises in pairs (a higher-order process), while gating or substrates can break mirror symmetry and activate linear ZA coupling.[94, 93, 95] In practice, lattice thermal conductivity in graphene is most sensitive to support/encapsulation, disorder and strain, with electrostatic control mainly influencing electronic transport. Dielectric environment and symmetry breaking can nevertheless feed back onto phonon lifetimes through changes in long-range fields and selection rules.[96]

### v2

@@ 4.2 The role of different phonon branches
In covalent semiconductors such as Si and Ge, long-wavelength acoustic (LA/TA) modes couple to carriers through the (short-range) deformation-potential interaction. The corresponding e–ph matrix elements vanish at $q\!\to\!0$ and increase with $|q|$; together with the linear acoustic dispersion $\omega=v_{s}q$, this makes carrier scattering by the very lowest-frequency LA/TA modes weak. As a result, these modes retain long mean free paths and dominate the lattice thermal conductivity at room temperature.[Agapito2018_PRB, Li2015_PRB, Broido2007_APL]
In polar III–V semiconductors the situation changes: the long-range Fröhlich field couples carriers strongly to longitudinal-optical (LO) phonons, and the mode dependence of the e–ph matrix elements becomes essential.[Verdi2015_PRL, Sjakste2007_PRL, Sjakste2015_PRB] When electrons and phonons are treated self-consistently in coupled BTE frameworks, momentum/energy exchange mediated by e–ph scattering can noticeably reduce the phonon thermal conductivity relative to a phonon–phonon-only model, especially at finite carrier densities.[Protik2022_npj, Protik2020_PRB]
Two-dimensional semiconductors add a quadratic flexural (ZA) branch. DFPT together with anharmonic renormalization shows that, at room temperature and for carrier densities $\lesssim 10^{12}\,\mathrm{cm^{-2}}$, ZA scattering contributes less than $\sim\!30\%$ of the total resistive e–ph rate; in-plane acoustic modes thus dominate both carrier mobility and lattice heat conduction under typical conditions.[Rudenko2019_PRB]
@@ 4.3 Carrier density and screening
where $D_{\mathrm{ac}}$ is the acoustic deformation potential, $\rho$ the mass density, $v_{s}$ the sound velocity, and $C\sim\mathcal{O}(1)$ is an angular factor that depends weakly on band anisotropy and degeneracy[Ziman1960, Li2015_PRB].
Second, the dense electron gas strengthens electrostatic screening. Rather than equating the screening vector to $k_{F}$, one should use a Thomas–Fermi/Lindhard wavevector $k_{s}$ that increases with $n$ (via $k_{F}$) but is not equal to it; transport coefficients are notably sensitive to the choice of $k_{s}$[Kukkonen1973_PRB]. In ab initio BTE studies of $n$-type Si, once $n\!\gtrsim\!10^{19}\,\mathrm{cm^{-3}}$ the calculated electron-limited damping of low-frequency LA/TA modes grows rapidly and the resulting suppression of $\kappa_{\mathrm{p}h}$ tracks time-domain thermoreflectance benchmarks[Li2015_PRB, Protik2022_elphbolt].
@@
so increasing $k_{s}$ (heavier doping) softens the long-range field. Nonetheless, room-temperature energy relaxation and phonon damping remain LO-dominated over wide doping ranges in nitrides; screening reduces the rate but does not eliminate the LO bottleneck[Jhalani2020_PRL, Zhang2001_JPCM]. In short, heavy doping in covalent Si throttles heat flow primarily by empowering low-frequency acoustics, whereas in polar GaN it leaves the LO channel as the primary sink—highlighting how bonding character (deformation potential vs. Fröhlich) and carrier screening jointly decide which branch ultimately limits the lattice thermal conductivity.
@@ 4.4 Dimensionality and dielectric environment
Reducing a semiconductor to monolayer or ultrathin‐film thickness fundamentally modifies the microscopic ingredients of the e–ph interaction. Confinement quantises the out‐of‐plane phonon wave vector, replacing the 3D continuum $q_{z}$ by discrete values $q_{z}=n\pi/L$ ($n=0,1,\dots$), while the in‐plane component $\mathbf{q}_{\parallel}$ remains continuous down to $\mathbf{q}_{\parallel}\!\to\!0$. The in‐plane acoustic branch with $n=0$ remains gapless with $\omega=v_{s}q_{\parallel}$; by contrast, the first thickness-quantised subband ($n=1$) acquires a gaplike term,
@@
so long‐wavelength 3D channels that require finite $q_{z}$ below $\pi/L$ are kinematically excluded even though arbitrarily small in‐plane momenta are allowed. A simple way to encode this thickness cutoff in the out‐of‐plane phase space is
@@
This $\Theta$-filter is a heuristic to encode the loss of continuous $q_{z}$ phase space; in real films the spectrum consists of discrete subbands (Lamb modes) rather than a hard cutoff. At the same time, freestanding monolayers host a quadratic flexural (ZA) branch with $\omega_{\mathrm{ZA}}=Aq^{2}$. For crystals with horizontal mirror symmetry, the leading e–ph coupling to ZA modes is symmetry‐forbidden at first-order and emerges at higher order, giving a small‐$q$ scaling $\lvert g_{\mathrm{ZA}}\rvert\!\propto\!q^{2}$; as a result, the intrinsic ZA contribution to resistive scattering is strongly suppressed compared to in‐plane acoustics unless symmetry is broken by substrates, strain, or gating.[Rudenko2019_interplay] First‐principles transport studies in 2D semiconductors consistently find that in‐plane acoustic modes dominate and that the role of flexural modes is reduced to $\lesssim\!30\%$ at low carrier densities (with further suppression on substrates).[Rudenko2019_interplay, Liao2015_Phosphorene]
Confinement also shortens relevant mean‐free paths compared with bulk. In monolayer MoS2, electronic mean free paths at 300 K are typically below $\sim\!9$ nm when phonon scattering limits transport[Li2015_PRB]. This has been shown in Fig. 3(a). phonon MFPs likewise contract relative to the bulk parents when boundary, substrate, and e–ph mechanisms are included, though precise values are material- and environment‐dependent.
A second confinement effect is electrostatic: the long‐range Fröhlich interaction in 2D depends on an effective out‐of‐plane permittivity set by the surrounding media,
@@
[…]
@@ 4.5 Coupled BTE and the role of elphbolt
The prevailing ab initio workflow treats electrons and phonons in parallel: one solves a phonon Boltzmann transport equation (BTE) for $\kappa_{\mathrm{p}h}$ and an electron BTE for $\sigma$ and $S$. In reality the two distribution functions are coupled: non-equilibrium electrons push phonons out of equilibrium (electron drag), and drifting phonons pull electrons along (phonon drag). Historically, semi-analytical drag models partially decoupled the BTEs and introduced effective terms, but this can violate the Kelvin–Onsager relationship connecting Seebeck and Peltier responses[Protik2022_npj, Sondheimer1956, Herring1954, Zhou2015_PNAS, Mahan2014_JAP].
@@
elphbolt solves the fully coupled, linearised electron–phonon BTEs from first principles. It iterates the electron and phonon collision integrals to self-consistency under the same driving fields ($\nabla T$ and $E$), and explicitly enforces the Kelvin–Onsager reciprocity at each iteration. Crucially, it uses exactly the same DFPT–Wannier electron–phonon matrix elements as single-species BTE workflows, so one can switch from the independent-channel picture to the coupled one without refitting any parameters[Giustino2017_RMP, Protik2022_npj, Sjakste2015_Wannier, Verdi2015_Frohlich].
@@
Concrete results illustrate when coupling matters. In $n$-type Si, elphbolt reproduces the temperature dependence of the Seebeck coefficient across low and high doping; the phonon-drag contribution becomes dominant below $\sim\!175$ K, while at room temperature direct e–ph scattering dominates and single-BTE and coupled-BTE predictions nearly coincide. For $\kappa_{\mathrm{p}h}$ in Si, the electron drag correction is small; suppression of $\kappa_{\mathrm{p}h}$ at low temperature is governed mainly by the phonon–electron scattering rates rather than mutual drag. Similar “small drag on $\kappa_{\mathrm{p}h}$” conclusions hold for GaAs and SiC[Protik2022_npj, Protik2020_GaAs, Protik2020_SiC]. This has been shown in Fig. 3(b). In strongly polar materials (e.g. GaN), accurate treatment of long-range polar couplings may require quadrupolar corrections; until those are included, drag conclusions should be drawn with care[Protik2022_npj, Sjakste2015_Wannier, Verdi2015_Frohlich].
@@
In short: use coupled e–ph BTEs when phonon or electron drag is relevant (low-to-intermediate $T$, clean samples, or structures with suppressed impurity scattering), or when enforcing Kelvin–Onsager consistency matters for quantitative thermoelectric predictions. Otherwise, single-BTE solutions (RTA or iterative) are often sufficient and substantially cheaper[Protik2022_npj].
@@
Figure 3: (a) In monolayer MoS2 at 300 K, phonon-limited electronic mean free paths cluster in the few–10 nm range (often <10<\!10 nm), illustrating the strong size/confinement sensitivity and providing context for the 2D transport lengths discussed in this subsection. Figure taken from Li et al.[Li2015_PRB]. (b) Lattice thermal conductivity of Si vs temperature at low and high doping from the fully coupled e–ph BTE agrees with experiment Inyushkin et al.[inyushkin2004isotope] and shows stronger low-$T$ suppression from e-ph scattering. At room temperature ph–ph scattering dominates and coupled/uncoupled predictions converge. Figure taken from Protik et al.[Protik2022_npj]
@@
A consistent body of measurements now benchmarks the fully first-principles workflow (DFPT/Wannier e–ph vertices + BTE solvers). In boron-doped silicon, thermal‐conductivity data on both natural and isotopically enriched crystals, together with doping-dependent measurements, exhibit the expected suppression of $\kappa_{\mathrm{p}h}$ as carriers are added and temperature is lowered. Fully coupled e–ph BTE calculations (elphbolt) reproduce the measured temperature dependence of $\kappa_{\mathrm{p}h}$ for nn‐type Si across dilute and degenerate limits, and quantify why the phonon–electron channel suppresses $\kappa_{\mathrm{p}h}$ more strongly at low $T$ (where long-wavelength acoustics dominate) than at high $T$ (where ph–ph scattering wins)[Protik2022_npj]. These comparisons are made directly against high-purity Si data and are consistent with isotope‐controlled measurements (which isolate mass-disorder effects) and with earlier doping-dependent thin-film studies[Inyushkin2018_SiIsotope, Asheghi2002_SiFilms]. [^4] In elphbolt’s Si case study, the measured points for $\kappa_{\mathrm{p}h}(T)$ at low and high doping (including $n\!\sim\!2\times 10^{19}\,$cm-3) are reproduced without empirical lifetimes, and electron-drag corrections to $\kappa_{\mathrm{p}h}$ are found to be small; see Fig. 6 and discussion in Ref. [Protik2022_npj].
@@
For polar III–V semiconductors (e.g. GaAs), experiments using transient thermoreflectance and related pump–probe probes report reductions of lattice thermal conductivity compatible with LO-phonon damping at finite carrier densities. First-principles treatments that include the long-range polar (Fröhlich) vertex capture these trends; coupled e–ph BTE solutions further show that, while phonon drag can dominate the Seebeck response at low $T$, the electron-drag correction to $\kappa_{\mathrm{p}h}$ remains small in GaAs, consistent with the mode-resolved balance of ph–e versus ph–ph scattering[Protik2022_npj, Sjakste2007_PRL].
@@
In 2D semiconductors, measurements on monolayer MoS2 consistently find a strong substrate/encapsulation dependence of the in-plane thermal conductivity, with reported room-temperature values for suspended or weakly coupled monolayers typically tens to $\sim\!$80 W m-1 K-1 and lower values on amorphous oxides[Zhu2023_RSC, Taube2015_MoS2_Raman]. These trends are aligned with first-principles expectations: the dielectric environment tunes the long-range polar coupling, and the dominant in-plane acoustic branches set the heat flux; ab initio transport in monolayer MoS2 also finds short electronic mean free paths and comparatively modest Fröhlich coupling in the intrinsic limit[Li2015_PRB]. Encapsulation in high-quality h-BN further improves interfacial heat dissipation and can raise the effective thermal conductance of hBN/MoS2/hBN stacks[Ye2021_hBNencap].
@@
Taken together—covalent Si (with isotope/doping controls), polar GaAs (LO-damping trends), and substrate-tunable 2D MoS2—these benchmarks show that once DFPT e–ph matrix elements are in hand, first-principles transport reproduces both magnitudes and temperature trends of $\kappa_{\mathrm{p}h}$ without empirical lifetimes, while clarifying where mutual e–ph drag matters and where it does not[Protik2022_npj].
@@ 4.6 Summary and open challenges
The next frontier is to widen this ab initio umbrella. First, correlated or magnetic semiconductors call for beyond-DFT band structures (DFT+U+U, hybrid functionals, or even GWGW/DMFT) so that the subsequent e–ph couplings rest on accurate quasiparticle states.[Giustino2017_RMP] Second, van der Waals heterostructures demand a truly nonlocal description of screening: the effective Fröhlich vertex in, e.g., MoS2/h-BN/graphene stacks depends on the full dielectric profile rather than a single scalar constant, and first-principles long-range corrections must be retained explicitly.[Verdi2015_PRL, Sjakste2015_Wannier] Third, pump–probe spectroscopy pushes materials into ultrafast, far-from-equilibrium regimes where hot carriers and nonthermal phonons coevolve; coupling the self-consistent electron– and phonon–BTE engine (elphbolt) to optical drive/relaxation protocols promises a parameter-free view of picosecond energy flow and drag phenomena.[Hartland2011_ChemRev, Protik2022_npj]
@@
Dirac crystals are gapless systems whose low–energy carriers follow a linear (Dirac) dispersion rather than the usual parabolic one. In graphene, conduction and valence cones touch at discrete Dirac points, and carriers near these points are well described by a massless Dirac Hamiltonian with an approximately constant Fermi velocity $v_{F}$.[CastroNeto2009_RMP, DasSarma2011_RMP] This “relativistic” band structure underlies characteristic transport phenomena such as the unconventional (half-integer) quantum Hall effect and Klein tunnelling.[CastroNeto2009_RMP, Mariani2008_PRL]
@@
In pristine, charge-neutral graphene the Fermi “surface” collapses to a point and the density of states vanishes linearly at the Dirac point. Modest electrostatic gating shifts $E_{\mathrm{F}}$ by tens to hundreds of meV, and—since the 2D Dirac DOS grows $\propto|E|$—yields large changes in carrier density.[CastroNeto2009_RMP, DasSarma2011_RMP] Doping then affects electron–phonon (e–ph) physics in two ways: it enlarges the electronic phase space for scattering and strengthens screening. The net impact is mode- and $\mathbf{q}$-selective: short-range deformation-potential couplings are efficiently screened, whereas the “gauge-field” part of the in-plane acoustic coupling remains essentially unscreened; moreover, in suspended graphene the linear coupling of out-of-plane (ZA) flexural modes is symmetry-forbidden at first-order and arises in pairs (a higher-order process), while gating or substrates can break mirror symmetry and activate linear ZA coupling.[SohierThesis2015, Mariani2008_PRL, Sohier2017_PRB] In practice, lattice thermal conductivity in graphene is most sensitive to support/encapsulation, disorder and strain, with electrostatic control mainly influencing electronic transport. Dielectric environment and symmetry breaking can nevertheless feed back onto phonon lifetimes through changes in long-range fields and selection rules.[NikaBalandin2017_RPP]
5.1 Electronic screening: Thomas–Fermi vs Lindhard (static/dynamic)